OpenStudy (anonymous):

a _{n}=((-1)^{n-1}(6)+2)/(2n+1+(-1)^{n-1}) for a positive interger k, find the simplified expression for \[a _{2k-1}\]

OpenStudy (anonymous):

\[a _{n}=((-1)^{n-1}(6)+2)/(2n+1+(-1)^{n-1})\]

OpenStudy (experimentx):

\[ a _{2k - 1}=((-1)^{2k-2}(6)+2)/(4k-1+(-1)^{2k-2}) \]

OpenStudy (anonymous):

how would I do \[a _{2k-1}+a _{2k}\]

OpenStudy (experimentx):

\[ a _{2k-1}=((-1)^{2k-2}(6)+2)/(4k-1+(-1)^{2k-2}) \] \[ a _{2k}= ((-1)^{2k-1}(6)+2)/(4k+1+(-1)^{2k-1}) \] \[ \frac{((-1)^{2k-2}(6)+2)}{(4k-1+(-1)^{2k-2})} + \frac{((-1)^{2k-1}(6)+2)}{(4k+1+(-1)^{2k-1})}\] \[ \frac{6+2}{4k-1+1} + \frac{-(6)+2}{(4k+1-1)}\]

OpenStudy (experimentx):

(-1)^2k-2 is +1 (-1)^2k-1 is -1

OpenStudy (anonymous):

are they not both over 4k and can be added easily?